OpenStudy (anonymous):

If f(x) = 5x - 9, find the value of f(-2). A. 8 B. -19 C. 23 D. -23

OpenStudy (anonymous):

\[f(x)=5x-9\] <=> \[f(-2)=(5\cdot-2)-9\] <=> \[f(-2)=\]

OpenStudy (anonymous):

I don't know how to do this lol

OpenStudy (phi):

replace x with -2 in the formula. erase the x and put in -2 do the arithmetic just as HomeZone shows

OpenStudy (phi):

you can multiply -2 times 5, right ?

OpenStudy (anonymous):

\[x=-2\] in your linear function \[f(x)=ax+b\]

OpenStudy (anonymous):

you insert \[x=-2\] into your function and find \[f(x)=y\]when \[x=-2\]=>\[f(-2)=(5\cdot-2)-9\]

OpenStudy (anonymous):

-10?

OpenStudy (anonymous):

to make it easy on you, you just gotta calculate: \[(5\cdot -2)-9=\] no it's not -10.

OpenStudy (anonymous):

Do you know P.E.M.D.A.S ?

OpenStudy (anonymous):

−19

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Yes, thats correct :)!

OpenStudy (phi):

I think musa did -2*5 to get -10 then the next step is subtract 9 to get -19 It looks like musa got the correct answer.

OpenStudy (anonymous):

ahh i see :P